OpenStudy (anonymous):

Please Help!! Can this formula x^2 - x +41 produce a composite number when x is substituted with a number?

OpenStudy (anonymous):

x=41

OpenStudy (anonymous):

Please show me, I have been getting all prime numbers

OpenStudy (anonymous):

plug in 41, your gonna get a composite number

OpenStudy (anonymous):

Thank You.. shoot I have been trying all numbers, haven't been in math in a long time and when you don't use it you lose it. : D

OpenStudy (anonymous):

True :), im getting dumber everyday

OpenStudy (anonymous):

41^2 - 41 + 41 = 1681 - 41 + 41 + =1681

OpenStudy (anonymous):

right, and 41 is a factor of 1681

OpenStudy (anonymous):

41 is a prime number and it when plugged into the spot it =1681, which can not be divided evenly.

OpenStudy (anonymous):

my understanding is that composite numbers are ones that can be divided evenly.

OpenStudy (anonymous):

I am getting prime numbers